OpenStudy (ilovecake):

Which is the best deal for a box of cereal A. $3.56 for 14 oz B. $9.13 for 39 oz C.$6.12 for 24 oz D.$4.98 for 18 oz

OpenStudy (doc.brown):

How much is one ounce of each?

OpenStudy (ilovecake):

I nee help!!

OpenStudy (doc.brown):

I'll do A, if you do the rest\[\dfrac{$3.56}{14oz}=\frac{$0.254}{1oz}\]

OpenStudy (ilovecake):

it is b I think

OpenStudy (doc.brown):

Why do you think so, is it because\[\frac{$9.13}{39oz}=\frac{$0.234}{oz}\]and that's less than $0.254/oz?

OpenStudy (ilovecake):

yes

OpenStudy (ilovecake):